# From Unimodal to Multimodal: Scaling up Projectors to Align Modalities

### v1

@@ Abstract
Recent contrastive multimodal vision-language models like CLIP have demonstrated robust open-world semantic understanding, becoming the standard image backbones for vision-language applications due to their aligned latent space. However, this practice has left powerful unimodal encoders for both vision and language underutilized in multimodal applications which raises a key question: Is there a plausible way to connect unimodal backbones for zero-shot vision-language tasks? To this end, we propose a novel approach that aligns vision and language modalities using only projection layers on pretrained, frozen unimodal encoders. Our method exploits the high semantic similarity between embedding spaces of well-trained vision and language models. It involves selecting semantically similar encoders in the latent space, curating a concept-rich dataset of image-caption pairs, and training simple MLP projectors. We evaluated our approach on 12 zero-shot classification datasets and 2 image-text retrieval datasets. Our best model, utilizing DINOv2 and All-Roberta-Large text encoder, achieves 76 accuracy on ImageNet with a 20-fold reduction in data and 65 fold reduction in compute requirements. The proposed framework enhances the accessibility of model development while enabling flexible adaptation across diverse scenarios, offering an efficient approach to building multimodal models by utilizing existing unimodal architectures. Code and datasets will be released soon.
@@ 1 Introduction
Contrastive multimodal vision-language models have demonstrated impressive zero-shot capabilities Radford et al. (2021); Jia et al. (2021); Zhai et al. (2023). These advancements have facilitated the use of language as an API for vision tasks, with captions functioning as adaptive classes, enabling a wide range of applications. The typical objective function InfoNCE, maximizes mutual information between the global summary vector of an image and its representation. However, the use of pooling functions to create global representations poses challenges for granular tasks like segmentation, which require pixel-level features. Notably, across various vision-centric benchmarks, unimodal models such as DINOv2 significantly outperform CLIP-like models Tong et al. (2024a; b). As the field progresses, there is an increasing demand for multimodal systems that can efficiently adapt to new modalities and tasks without extensive retraining. This evolution highlights the need for more flexible and efficient approaches to multimodal learning.
While efforts have been made to develop more efficient CLIP-like models, they often compromise on performance or still demand significant resources. For instance, LiT Zhai et al. (2022) achieves comparable performance to CLIP but still requires training on 256 TPU cores with over 4 billion image-caption pairs. Smaller-scale attempts like LiLT Khan & Fu (2023) show promise in retrieval tasks but struggle with zero-shot classification accuracy.
In this paper, we propose an alternative approach to vision-language multimodal alignment that strives to address these challenges. Our method builds upon the recent findings suggesting semantic similarities between well-trained unimodal vision and language embeddings Maniparambil et al. (2024); Huh et al. (2024). By leveraging these semantic similarities, we explore the potential of creating efficient CLIP-like models by training lightweight projection layers between unimodal frozen models.
@@
This approach has two practical benefits compared to CLIP-like models:
Accessible development: Training only projection heads with a dense dataset significantly reduces the computational requirements compared to full model training. This approach not only decreases the environmental impact of developing foundation models but also makes their creation more accessible to the broader research community (see Section 5.5 for detailed comparisons).
Flexible adaptation to diverse scenarios: By connecting unimodal encoders through lightweight projectors, our method enables the utilization of specific features from each encoder in a multimodal context. Examples include: (1) Creating multilingual vision-language models for low-resource languages by aligning DINOv2 with a multilingual text encoder using a small set of image-caption pairs in the target language (see Section 5.3), as well as (2) Enabling image-paragraph retrieval by aligning visual encoders with long-context language models like BERT-large, overcoming the token limit constraints of standard CLIP models (See Section 5.4).
Our framework consists of three key components:
1.
Encoder Pair Selection: We identify semantically similar vision and language encoders using the Centered Kernel Alignment (CKA) metric.
2.
Dataset Curation: We develop a method to collect a dense, concept-rich dataset of image-caption pairs from uncurated sources. We argue that alignment is sensitive to concept coverage, and carefully select samples that cover most of the target concepts.
3.
Lightweight Projector Training: We train simple MLP projectors between the embedding spaces of frozen unimodal models using contrastive loss.
@@
We evaluate our approach on zero-shot transfer to 12 different classification datasets and 2 image-text retrieval datasets. Our best projector between unimodal models, utilizing DINOv2 and All-Roberta-Large-v1, achieves 76% accuracy on ImageNet, surpassing CLIP’s performance while using approximately 20 times less data and 65 times less compute. We also demonstrate our framework’s versatility across tasks like zero-shot domain transfer, multilingual classification, zero-shot semantic segmentation, and image-paragraph retrieval.
@@
2 Selecting which encoders to align
@@
Previous studies Huh et al. (2024); Maniparambil et al. (2024) have shown that well-trained vision and language encoders exhibit high semantic similarity, as measured by Centered Kernel Alignment (CKA) Kornblith et al. (2019) (see Section A.1 for CKA details). A layerwise analysis in Maniparambil et al. (2024) reveals that most of this similarity is concentrated in the final projection layer. Additionally, model stitching methods Lenc & Vedaldi (2015); Bansal et al. (2021); Merullo et al. (2022) demonstrate that different network regions can be stitched together using linear layers. Inspired by this, in Section 2.1, we investigate whether semantically similar encoder embedding spaces can be aligned through a simple projection transformation, using toy examples in the first instance to validate the underlying concept.
2.1 CKA vs Ease of Alignment
@@
Centered Kernel Alignment (CKA) Kornblith et al. (2019) measures the similarity of induced graphs of concepts in each encoder space and can act as a guide for selecting encoder pairs that are amenable to alignment. Here we aim to understand the relationship between CKA and the level of difficulty of aligning embedding spaces using toy examples. We define the ease of alignment as the minima of the training loss after convergence. We examine how CKA correlates with the minimum CLIP loss when transforming one vector set to match another using a Linear layer. Since CLIP loss lacks a closed-form solution, we applied SGD for iterations per instance, recording the final loss value as the minimum. We fixed the temperature at and the learning rate at , choosing iterations because the loss value showed minimal change beyond this point.
In our experiment, we generated instances of two vector sets, and , each containing 32 vectors of 16 dimensions. Following the approach in Maniparambil et al. (2024); Huh et al. (2024), we modeled the world using a latent distribution , with Image and Text representations ( and ) as random independent non-linear transformations from with additive noise. For each sampled pair of and matrices, we calculated the CKA and the minimum CLIP loss. The non-linear transform was defined as a randomly initialized 2-layer MLP with ReLU non-linearity and hidden dimensions significantly larger than the input dimensions, ensuring it could universally approximate the non-linear transformation Hornik et al. (1989). Figure 2 was used to generate each instance.
Figure 1(a) illustrates the results of this experiment, showing a clear negative correlation between CKA and minima of the CLIP loss. As CKA increases, indicating greater similarity between the similarity structures of A and B, the minima of CLIP loss consistently decreases. Despite arising from a simplified experiment, this strong inverse relationship provides empirical support for the use of CKA as a predictor of alignment potential between different embedding spaces.
2.2 Ease of Alignment with real embeddings
@@
We go a step further and investigate whether a similar inverse relationship exists between the minima of CLIP Loss and CKA when the embeddings are from real data and real encoders. We consider 35 different sentence encoders and 25 different vision encoders and sample 3000 different image,caption pairs from the COCO validation set and pass them through all possible encoder combinations to produce 600 different sets of A and B. We then calculate CKA and compute the minima of CLIP Loss after 500 iterations for these A, B and plot them in Figure 1(b) with CKA on the x axis and minima of CLIP loss on the y axis in a log scale. We see that even for real-world embeddings of images and captions there is a strong inverse relationship between CKA and the minima of the CLIP loss providing further evidence that encoders with high CKA could have similar similarity structures making them easy to align using simple projections.
@@ 3.1 Encoder Pair Selection
Inspired by Section 2 we use CKA for selecting the most semantically similar encoder pairs for multimodal alignment. We opted for a linear kernel in the CKA computation after observing that the trends in results were largely consistent between linear and RBF kernels, while the linear kernel offers superior computational efficiency. We measure the CKA between encoder spaces by constructing sets of vision embeddings and text embeddings on the COCO validation set of 5000 image, caption pairs. The COCO validation set is chosen as the reference set for its high semantic alignment between the image content and the caption description. We ablate the use of CKA for encoder pair selection in 4.1 and find a positive correlation between CKA and transfer performance to downstream datasets.
@@ 3.2 Dataset Curation
By only training the projection layers to align embedding spaces, our approach requires significantly less data compared to training a CLIP model from scratch. To achieve high-quality alignment, it is essential to use a small but well-curated dataset featuring image-caption pairs with a strong semantic correspondence between the images and the text. Additionally, the dataset must encompass a wide range of concepts to facilitate robust zero-shot domain transfer capabilities. With these requirements in mind, our dataset curation process is structured into three key steps:
Concept Image Prototypes: Firstly, we collect 3000 unique concepts from class names of ImageNet, and several other datasets (see A.10.1). Then, we embedded 128 image samples for each concept from the corresponding dataset using CLIP VIT-Large’s vision encoder and average them to obtain the concept image prototypes.
Concept Level Collection: To create a class-balanced dataset, we first collect image-caption pairs from LAION400M, a large, uncurated source dataset. We then embed all captions using CLIP ViT-Large’s text encoder and compute the caption-image prototype similarity for each concept. To ensure diversity, we retrieve 2,000 samples per concept, starting with the less common concepts. As a proxy to establish the commonality of a concept in the pool, we use the average cosine similarity of the top 25,000 captions closest to each concept prototype. This process results in LAION-CLASS-Collected, a high-quality dataset of 6M samples with broad concept coverage. The detailed algorithm is illustrated in Fig 3, and A.5 details the implementation and compute requirements for our collection process.
Retrieval Datasets: The LAION-CLASS-Collected dataset offers high concept diversity, but LAION itself is uncurated, with many captions poorly aligned with their images Fan et al. (2024); Nguyen et al. (2024); Chen et al. (2023b). While concept coverage is crucial for strong zero-shot classification, image quality, text diversity, and image-caption alignment are key for effective zero-shot image-text retrieval. In contrast, datasets like CC3M Sharma et al. (2018), CC12M Changpinyo et al. (2021), and SBU Ordonez et al. (2011) feature higher-quality images and better image-caption alignment than LAION. By combining these, we create a 20M MIX-CLASS-Collected dataset that balances concept coverage with image-text similarity, enhancing both retrieval performance and zero-shot domain transfer across various classification tasks. We examine the impact of each data source on task performance in Sec 4.3.
@@ 3.3 Projector Architecture
We train lightweight projectors using contrastive loss between adapted image and text embeddings while keeping the unimodal encoders frozen. Figure 4 shows our projector architecture/configuration. We use a lightweight Token Projector Mukhoti et al. (2023) with linear and non-linear branches in a residual configuration for both local tokens and the CLS token of each encoder. The projector’s weights are shared for local tokens and separate for the CLS token. Adapted local tokens are averaged and added to the adapted CLS token to form a global embedding, capturing both global and local encoder information.
For sentence-transformer architectures, Token Projectors are applied to the tokens, followed by a 2-layer MLP as a global Text Projector, as the text embeddings need further adaptation to become more aligned with the vision embeddings. All projector choices are thoroughly ablated in Section 4.2. Training information and hyperparameters are detailed in A.6.
4 Ablation Experiments
We present a set of ablations to validate different components of our pipeline empirically: CKA for encoder selection 4.1, the projector architecture and configuration 4.2, the alignment datasets, and the impact of class-collected data 4.3. We evaluate on downstream tasks like 0-shot domain transfer to Imagenet classification and COCO / Flickr30k image-text retrieval scores.
4.1 Effectiveness of CKA for encoder pair selection
We train our projector configurations on various combinations of unimodal encoders using the COCO dataset and evaluate image/text retrieval accuracies on the Flickr30k test set, plotting these against CKA scores in Figure 6. The CKA, calculated on the COCO image-caption pairs, shows a strong correlation with retrieval accuracy, indicating that higher semantic similarity, as measured by CKA, predicts better alignment in image/text retrieval. Our findings suggest that CKA can effectively predict which encoder pairs will align well with projector training. The DINOv2-Large and CLIP-ViT-Large-text combination achieves the highest retrieval score, but certain unimodal pairs, like DINOv2-Large and All-Roberta-Large-v1 (CKA = 0.69), perform nearly as well. This indicates that these unimodal encoders are highly effective for vision-language alignment, leading us to choose the DINOv2-Large and All-Roberta-Large-v1 pair for larger-scale experiments. Image Retrieval performance is illustrated in A.2.
Unimodal Performance Does Not Reflect Alignment Quality: We perform the same ablation as above using DINOv2 and 14 different text encoders from the Sentence Transformers library Reimers & Gurevych (2019). In Fig. 5, we plot Flickr30k text retrieval accuracies against text encoder performance averaged over sentence embedding (STS) tasks (14 datasets) and semantic search (SS) tasks (6 datasets). The results show that text encoder performance does not predict alignment quality, suggesting that CKA, rather than unimodal performance, can be used to identify encoder pairs that easily align. Further ablations are discussed in A.4.

### v2

# Harnessing Frozen Unimodal Encoders for Flexible Multimodal Alignment

@@ Abstract
Recent contrastive multimodal vision-language models like CLIP have demonstrated robust open-world semantic understanding, becoming the standard image backbones for vision-language applications. However, recent findings suggest high semantic similarity between well-trained unimodal encoders, which raises a key question: Is there a plausible way to connect unimodal backbones for vision-language tasks? To this end, we propose a novel framework that aligns vision and language using frozen unimodal encoders. It involves selecting semantically similar encoders in the latent space, curating a concept-rich dataset of image-caption pairs, and training simple MLP projectors. We evaluated our approach on 12 zero-shot classification datasets and 2 image-text retrieval datasets. Our best model, utilizing DINOv2 and All-Roberta-Large text encoder, achieves 76 accuracy on ImageNet with a 20-fold reduction in data and 65-fold reduction in compute requirements compared multi-modal alignment where models are trained from scratch. The proposed framework enhances the accessibility of multimodal model development while enabling flexible adaptation across diverse scenarios. Code and curated datasets are available at github.com/mayug/freeze-align.
@@ 1 Introduction
Contrastive multimodal vision-language models have recently demonstrated impressive zero-shot capabilities [45, 22, 62]. These advancements facilitate the use of language as an API for vision tasks, treating captions as adaptive classes to support a wide range of applications. However, current models face significant challenges: the typical objective function, InfoNCE, is designed to maximize mutual information between the global summary vector of an image and its text representation. This global approach, which relies on pooling functions within the CLIP vision encoder, struggles to deliver the pixel-level granularity required for tasks like segmentation [5]. In contrast, recent advances in uni-modal vision encoders, such as the DINOv2 [41], have demonstrated strong performance in both global and local vision tasks. The CLIP text encoder is limited by its English-only tokenizer and a fixed token length of 77, restricting it’s long-context and multilingual retrieval capabilities. Meanwhile unimodal language encoders [48], excel in multilingual, and long-context abilities, as evidenced by improved performance on MTEB benchmarks [36]. Despite these advances in unimodal models, the current strategy for aligning vision and language models usually involves full retraining of vision and language encoders, which is both computationally expensive and inflexible.
This paper proposes a framework for vision-language alignment that efficiently leverages advanced uni-modal vision and language encoders, creating adaptable multimodal models by training only projectors between their frozen embedding spaces. Current efforts to create more efficient CLIP models often compromise on either performance or still require significant resources. For example, LiT [63] achieves comparable results to CLIP but relies on massive compute resources, while smaller-scale models like LiLT [23] may lack sufficient concepts in their training datasets, limiting their zero-shot domain transfer accuracy.
To address these challenges, our approach builds on recent findings suggesting semantic similarities between well-trained unimodal vision and language embedding spaces [32, 21]. We hypothesize that these similarities enable effective alignment through simple projection transformations, and verify through a toy example in Section 3.2 and extensive ablation studies in Section 5.1. Inspired by this, our framework includes three key steps: identifying semantically similar vision-language encoder pairs, curating concept-dense datasets, and training lightweight projectors for efficient alignment.
@@
This approach has three practical benefits compared to CLIP-like training:
Strong Unimodal Features lead to Strong Multi-Modal Models Features from uni-modal vision and text encoders are more general than multi-modal trained encoders. For example, it’s been shown that vision-only trained encoders perform better on vision-centric tasks when compared to multi-modal vision encoders like CLIP-vision [55]. Hence by keeping these uni-modal encoders frozen and training only projectors for alignment, we aim to keep these strong uni-modal features intact, resulting in better multi-modal representations(See Sec. 6.2). Flexible adaptation to diverse scenarios: By utilizing the frozen unimodal encoders ability to handle a specific type of data we can efficiently train multimodal models that also can handle this specialized data without the need to retrain the whole network from scratch. For example, multilingual or long context vision-language models can be achieved by aligning DINOv2 with a multilingual (Section 6.3) or long-context language text encoder(Section 6.4). Accessible development and Model Reuse: Relying on already established encoders, projection heads with a dense dataset require significantly less computational resources compared to full model training. In purely practical sense, this approach not only decreases the environmental impact of developing multimodal models but also makes their creation more accessible to the broader research community (Section 6.5).
@@
Finally, we evaluate our approach on zero-shot transfer to 12 different classification datasets and 2 image-text retrieval datasets. Our best projector between unimodal models, utilizing DINOv2 and All-Roberta-Large-v1, achieves 76% accuracy on ImageNet, surpassing CLIP’s performance while using approximately 20 times less data and 65 times less compute for alignment. We also demonstrate our framework’s versatility across tasks like zero-shot domain transfer, multilingual classification, zero-shot semantic segmentation, and image-paragraph retrieval.
@@
2 Related Works
Multimodal Pretraining: The CLIP models from OpenAI [45] and ALIGN [22] pioneered using web-scale image-caption data to align image and text modalities via an InfoNCE [40] loss, optimizing mutual information between embeddings. LAION [51, 50] replicated this approach in the open domain, open-sourcing pre-training datasets. While these models excel in zero-shot tasks, they demand substantial computational resources, around 20k GPU hours. Taking advantage of the recent improvements in the representation quality of unimodal encoders such as DINOv2 [41] (vision) and Sentence Transformer [47] (language) models, [63] reduce the training cost by locking the image encoder and training only the text encoder to achieve competitive performance. Similarly, [23] further aligned frozen uni-modal encoders using projection layers, BitFit [61], and trainable adapters, but their approach is sub-optimal compared to CLIP, likely due to smaller datasets used and random encoder pair selection. In contrast, we strive to identify the best encoder pairs for alignment first and then scale up projector-only training to improve the multimodal alignment.
Representational Similarity: Recent studies show that the semantic similarity between vision and language model embeddings is high for several model pairs. [32] reports that this similarity, measured by Centered Kernel Alignment [24], increases with more training data for vision models. Similarly, [21] finds that better-performing language models have higher semantic similarity to the DINOv2 [41] vision model. These similarities have been leveraged for 0-shot and multi-lingual retrieval tasks using strong uni-modal encoders without additional training [32, 35], though scalability is an issue. Additionally, [34] demonstrates that a simple linear mapping allows a frozen language model to interpret visual input, provided the visual encoder aligns with language concepts (e.g., CLIP). These findings suggest that a simple projection transformation separates the embedding spaces of well-trained vision and language models, motivating our work on developing CLIP models using projection layers between semantically similar encoder pairs.
Automatic Data Curation: Our dataset curation pipeline draws on various approaches in Vision-Language dataset construction [45, 16, 60]. [45] used image metadata to gather high-quality image-caption pairs, while [51] replicated the CLIP dataset by filtering with pretrained vision encoders. Recent methods like [16] employ CLIP-based filtering and ad hoc filtering techniques, and [60] mimics CLIP’s data collection via metadata retrieval. Similarly, [41] uses a pretrained vision encoder to curate web images most similar to images in curated datasets. Our approach is similar, constructing concept image prototypes from few-shot labeled examples and retrieving relevant web images from the LAION-400M pool using CLIP caption embeddings, avoiding the computational cost of generating vision embeddings for the entire dataset.
3 CKA vs Ease of Alignment
@@
Previous studies [21, 32] have shown that well-trained vision and language encoders exhibit high semantic similarity using metrics like Centered Kernel Alignment. Specifically, a layerwise analysis in [32] reveals that most of this similarity is concentrated in the final projection layer. Furthermore, model stitching methods [26, 3, 34] demonstrate that different network regions can be stitched together using linear layers suggesting that deep representations that contain high-level semantics can be connected by simple transformations. Inspired by this, we investigate whether semantically similar embedding spaces can be aligned through a simple projection transformation, using a toy example to validate the underlying concept.
3.1 CKA Preliminary
Centered Kernel Alignment (CKA) has shown its relevance in understanding and comparing the information encoded by different layers of a neural network. CKA can be defined as follows: Given two sets of vectors and , CKA measures the similarity of these vectors in their respective high-dimensional feature spaces. The kernel matrices and are derived from the data sets and , respectively, and represent the inner products between the vectors in these spaces. The entries of and are:
The CKA coefficient, , is defined as:
where HSIC stands for Hilbert-Schmidt Independence Criterion [18, 30], which measures the dependence between the sets of vectors. This measure is invariant to orthogonal transformations and isotropic scaling of the data, making it robust for comparing different models.
3.2 CKA and Ease of Alignment Toy Example
@@
We define the Ease of Alignment as the minimum training loss achieved after convergence, reflecting the efficiency of aligning encoder outputs. We explore how Centered Kernel Alignment (CKA) correlates with the minimum CLIP loss when transforming one vector set to match another using a Linear layer. Given the lack of a closed-form solution for CLIP loss, we employ Stochastic Gradient Descent (SGD) for iterations per instance, recording the final loss as the minimum. We fixed the temperature at and the learning rate at , selecting iterations as the loss plateaued beyond this point.
@@
In this experiment, we examine if there is an inverse relationship between the minima of CLIP Loss and CKA for embeddings derived from real data using different language and vision encoders. We sample image-caption pairs from the COCO validation set and process them through five different sentence encoders and nine vision encoders, generating unique sets of embeddings (A and B). We calculate CKA and record the CLIP Loss after iterations for each set, plotting these values in Figure 1 with CKA on the x-axis and minima of CLIP loss on the y-axis on a log scale. The results confirm a strong inverse relationship between CKA and the minima of CLIP loss, suggesting that high CKA scores indicate similar structural similarities in encoders, which facilitate their alignment through simple projection methods. Further details on toy examples and visualization of similarity structures can be found in Sections A.4 and A.5.
@@ 4.1 Encoder Pair Selection
Inspired by Section 3 we use CKA for selecting the most semantically similar encoder pairs for multimodal alignment. We opted for a linear kernel in the CKA computation after observing that the trends in results were largely consistent between linear and RBF kernels, while the linear kernel offers superior computational efficiency. We measure the CKA between encoder spaces by constructing sets of vision embeddings and text embeddings on the COCO validation set of 5000 image, caption pairs. The COCO validation set is chosen as the reference set for its high semantic alignment between the image content and the caption description. We ablate the use of CKA for encoder pair selection in 5.1 and find a positive correlation between CKA and transfer performance to downstream datasets.
@@ 4.2 Dataset Curation
By only training the projection layers (11M parameters) to align embedding spaces, our approach requires significantly less data compared to training a CLIP model from scratch. However, to ensure high-quality alignment and effective transfer to diverse downstream tasks, it is essential to use a small but well-curated dataset that has the following features. 1. high concept coverage which aids in covering all regions of the uni-modal embedding spaces 2. high semantic alignment between image-caption pairs which aids in learning an effective mapping between vision and the embedding spaces. With these requirements in mind, our dataset curation process is structured into two key steps:
Concept Coverage Collection: To ensure high concept coverage, we collect 3000 unique concepts from class names of ImageNet, and several other curated datasets (see A.14.1). Concept image prototypes are then constructed by averaging few-shot image embeddings for each concept using CLIP VIT-Large’s vision encoder. To create a class-balanced dataset, we first collect image-caption pairs from LAION400M, a large, uncurated source dataset. We then embed all captions using CLIP ViT-Large’s text encoder and compute the caption-image prototype similarity for each concept. To ensure diversity, we retrieve 2,000 samples per concept, starting with the less common concepts. As a proxy to establish the commonality of a concept in the pool, we use the average cosine similarity of the top 25,000 captions closest to each concept prototype. This process results in LAION-CLASS-Collected, a high-quality dataset of 6M samples with broad concept coverage. The detailed algorithm is illustrated in Fig 2. A.8 details the implementation and compute requirements for our collection process.
Our primary goal is to compile a concept-rich dataset that enables quick learning and validates the efficacy of projectors for modality alignment, rather than developing a specific curation method. This paper demonstrates the potential of such multimodal models, emphasizing their practicality and efficiency when supported by a dataset with ample concept coverage and robust semantic alignment. The development of an exhaustive dataset that spans all domains of unimodal spaces, ensuring optimal semantic alignment between images and captions, is reserved for future work.
Retrieval Datasets: The LAION-CLASS-Collected dataset offers high concept diversity, but LAION itself is uncurated, with many captions poorly aligned with their images [15, 38, 10]. While concept coverage is crucial for a dense coverage of the unimodal embedding space, image quality, text diversity, and image-caption alignment are key for effective zero-shot image-text retrieval. In contrast, datasets like CC3M [52], CC12M [8], and SBU [42] feature higher-quality images and better image-caption alignment than LAION. By combining these, we create a 20M MIX-CLASS-Collected dataset that balances concept coverage with image-text similarity, resulting in both dense coverage of the uni-modal embedding spaces as well as high semantic alignment between cross-modal embeddings. We examine the impact of each data source on task performances in Sec 5.3.
@@ 4.3 Projector Architecture
We train lightweight projectors using contrastive loss between adapted image and text embeddings while keeping the unimodal encoders frozen. Figure 3 shows our projector architecture/configuration. We use a lightweight Token Projector [37] with linear and non-linear branches in a residual configuration for both local tokens and the CLS token of each encoder. The projector’s weights are shared for local tokens and separate for the CLS token to enable adaptation of both spatial and global information of the vision encoder while limiting the parameter count. Adapted local tokens are averaged and added to the adapted CLS token to form a global embedding, capturing both global and local encoder information. For text encoders, Token Projectors are applied to the tokens, followed by a 2-layer MLP as a global Text Projector, as the text embeddings need further adaptation to become more aligned with the vision embeddings. All projector choices are thoroughly ablated in Section 5.2. Training information and hyperparameters are detailed inA.9.
@@
5 Ablation Experiments
We present a set of ablations to validate different components of our pipeline empirically: CKA for encoder selection 5.1, the projector architecture and configuration 5.2, the alignment datasets, and the impact of class-collected data 5.3. We evaluate on downstream tasks like 0-shot domain transfer to Imagenet classification and COCO / Flickr30k image-text retrieval scores.
5.1 Effectiveness of CKA for encoder pair selection
We train our projector configurations on various combinations of unimodal encoders using the COCO dataset and evaluate image/text retrieval accuracies on the Flickr30k test set, plotting these against CKA scores in Figure 4. The CKA, calculated on the COCO image-caption pairs, shows a strong correlation with retrieval accuracy, indicating that higher semantic similarity, as measured by CKA, predicts better alignment in image/text retrieval. Our findings suggest that CKA can effectively predict which encoder pairs will align well with projector training. The DINOv2-Large and CLIP-ViT-Large-text combination achieves the highest retrieval score, but certain unimodal pairs, like DINOv2-Large and All-Roberta-Large-v1 (CKA = 0.69), perform nearly as well. This indicates that these unimodal encoders are highly effective for vision-language alignment, leading us to choose the DINOv2-Large and All-Roberta-Large-v1 pair for larger-scale experiments. Image Retrieval performance is illustrated in A.5. Additionally, our findings indicate that CKA serves as a more reliable and straightforward metric for assessing alignment quality compared to other encoder pair selection strategies, such as downstream task performances, which tend to vary significantly depending on the specific task chosen(See Sec A.7).